how do I solve (2y-9)+(y+4)=5y..what are the steps to solve these types of equations?
Answer by ankor@dixie-net.com(22740)

how do I solve (2y-9)+(y+4)=5y
There is a positive (+) between the brackets so we can just remove them
2y - 9 + y + 4 = 5y
combine like terms on the left
2y + y - 9 + 4 = 5y
3y - 5 = 5y
We want y on the left and numerical values on the right
Add 5 to both sides
3y = 5y + 5
Subtract 5y from both sides
3y - 5y = 5
-2y = 5
y has to be positive, multiply both sides by -1 to change the signs
2y = -5
we want y to have a coefficient of 1, divide both sides by 2
y = -5/2
y = -2.5
:
:
See if that checks out in the original equation, replace y with - 2.5
2(-2.5) - 9 - 2.5 + 4 = 5(-2.5)
-5 - 9 - 2.5 + 4 = -12.5
-12.5 = -12.5 confirms our solution of y = -2.5